# HOPE: A Reinforcement Learning-based Hybrid Policy Path Planner for Diverse Parking Scenarios

### v2

@@ Abstract
Automated parking stands as a highly anticipated application of autonomous driving technology. However, existing path planning methodologies fall short of addressing this need due to their incapability to handle the diverse and complex parking scenarios in reality. While non-learning methods provide reliable planning results, they are vulnerable to intricate occasions, whereas learning-based ones are good at exploration but unstable in converging to feasible solutions. To leverage the strengths of both approaches, we introduce Hybrid pOlicy Path plannEr (HOPE). This novel solution integrates a reinforcement learning agent with Reeds-Shepp curves, enabling effective planning across diverse scenarios. HOPE guides the exploration of the reinforcement learning agent by applying an action mask mechanism and employs a transformer to integrate the perceived environmental information with the mask. To facilitate the training and evaluation of the proposed planner, we propose a criterion for categorizing the difficulty level of parking scenarios based on space and obstacle distribution. Experimental results demonstrate that our approach outperforms typical rule-based algorithms and traditional reinforcement learning methods, showing higher planning success rates and generalization across various scenarios. We also conduct real-world experiments to verify the practicability of HOPE. The code for our solution will be openly available on https://github.com/jiamiya/HOPE.
@@ I Introduction
Automated parking is a tempting technology to improve driving safety and efficiency [1]. An automated parking system comprises several vital components, including perception, planning, and control, in which path-planning algorithms play a crucial role [2]. In parking scenarios, the path-planning task involves generating a feasible path from the start position to the target parking spots under specific physics constraints. Compared to other scenarios, finding a feasible path in parking scenarios is usually more challenging because of the lower error tolerance of the target spot and the lack of navigational reference lines [3]. Additionally, the limited space and surrounding obstacles shrink the number of potential solutions. While existing path-planning approaches have proved practical and robust in most simple scenarios, their inherent difficulty in truly understanding the surrounding environment may lead to planning failures, especially when scenarios become more intricate [4].
Learning-based planners have the potential to understand environments and intelligently plan routes through a data-driven approach, diverging from reliance on pre-defined planning methods rooted in human priors. While expert data is used as the ground truth for imitation learning in various tasks, the scarcity of large-scale datasets for parking scenarios necessitates researchers to collect data manually [5]. This supervised learning approach risks overfitting the model to specific parking strategies due to the insufficient diversity in the training scenarios. Meanwhile, Reinforcement Learning (RL) has gathered increasing attention in the field of autonomous driving [6]. Indeed, through interaction with the environment, RL methods enable the training of agents without labeling trajectory ground truth. Nonetheless, training RL agents in complex and diverse scenarios remains a challenging task [7]. It is straightforward for the agent to fit into fixed parking strategies but poses greater difficulty in obtaining generalization capability across various parking scenarios. Moreover, the agent faces challenges in conducting effective exploration, especially in complex scenarios with narrow parking spaces, significantly impacting the training efficiency.
@@
  • We develop a hybrid policy method for parking path planning which achieves over 97% success rate across diverse and challenging parking scenarios and verify its generalization ability with real-world experiments.
  • We propose a method for calculating and implementing the action mask in the path planning task. This mechanism excludes improper actions for reinforcement learning agents and significantly improves training efficiency and performance.
@@ III-A Reinforcement learning for path planning
The RL problem can be addressed as policy learning in a Markov decision process (MDP) defined by a 4-element tuple $(\mathcal{S},\mathcal{A},p,r)$, where $\mathcal{S}$ is the state space, and $\mathcal{A}$ is the action space. In the path-planning problem, $s_{t}\in\mathcal{S}$ includes the vehicle position and orientation $p_{t}=(x_{t},y_{t},\theta_{t})$ as well as other observable information about obstacles and target parking space. We selecte velocity $v$ and steering angle $\delta$ as the action space $\mathcal{A}=\{a=(v,\delta)\}$. $p$ denotes the state transition probability density of the next state $s_{t+1}$ given the current state $s_{t}$ and action $a_{t}$. In practice, the transition is modeled using the single-track bicycle model, as implemented in [32], and this uncertainty is not considered in a deterministic environment. A reward $r=r(s_{t+1},a_{t})$ is given by the environment $E$ based on the state and action in each interaction step. The objective is to learn the policy $\pi(a_{t}|s_{t})$ that maximizes the future reward $R_{t}=\sum_{i=t}^{T}\gamma^{(i-t)}r(s_{i},a_{i})$ with a discounting factor $\gamma\in[0,1]$. Given the start point $p_{0}$ and the target $p_{T}$, the feasible parking path $P=\{p_{0},p_{1},p_{2},...,p_{T}\}$ can be obtained iteratively using the policy $\pi$.
@@
To demonstrate the improvement of our proposed method across different reinforcement learning approaches, in this paper, we chose the commonly used on-policy algorithm, Proximal Policy Gridient (PPO) [33], and off-policy algorithm, Soft Actor-Critic (SAC) [34], to obtain the reinforcement learning policy $\pi_{\theta}$ in HOPE.
@@ IV-A Overview of architecture
The overall RL-based path planning framework and network structure are shown in Figure 1. To enhance training efficiency, we employ a hybrid policy reinforcement learning approach. Our proposed HOPE combines the learnable policy from original reinforcement learning with a rule-based policy derived from the RS curve. In each interaction step, the agent outputs the action, namely the single-step path planning result, based on the current state given by the environment. The action is then adjusted using the action mask mechanism before interacting with the environment. We employ four forms of network input as state representation, including information about the distance to obstacles, target parking space, drivable action space, and historical trajectory, encoded in either vectors or images. A transformer-based structure with learnable view encoding is designed to fuse the inputs and get the outputs in the actor and critic network [35]. We also utilize the auto-encoder structure to pre-train the image encoder. More details of our input modality and reward function are discussed in the appendix A and C. The pseudo-code of our method is shown in Algorithm 1.
@@ IV-B1 Rule-based policy from Reeds-Shepp curve
  • We calculated the shortest $K$ curves and verified their feasibility in ascending order of length. The shortest collision-free path is selected to calculate the planning action using the vehicle model.
  • We add the computation of the paths in expression $S(|)C(|)S$ as these paths, though strictly suboptimal in the absence of obstacles, may be feasible when obstacles are considered. Experiments in appendix B suggest this is effective for vertical parking.
@@ IV-C1 Efficient estimation of action mask
While solving for the optimal $l_{ij}$ considering only obstacle point at $i^{th}$ angle $\omega_{i}$ and vehicle steering angle $\delta_{j}$ is equivalent to obtaining the maximum velocity, denoted as $v^{*}_{ij}=\mathcal{E}_{ij}^{-1}(l_{t}[i])$, calculating all envelope distances corresponding to each $i$ and $j$ in each interaction is computationally expensive, and the inverse of $\mathcal{E}_{ij}$ may not always exist. To deal with this issue, we propose the pre-calculation of the anchoring distance $\hat{l}_{ij}$ for $K$ discretized velocities:
$\hat{l}_{ij}[k]=\mathcal{E}_{ij}(\hat{v}[k]),\ \hat{v}[k]=v_{max}\frac{k}{K},\ k=0,1,...K.$ (12)
@@ V-A Scenario difficulty ranking
To better train and evaluate our approach, we rank the difficulty for parking scenarios into normal, complex, and extreme levels based on existing standards ISO 20900 and GB/T 41630-2022 [37, 38]. Generally, a parking space is defined by two boundary parking vehicles (or boundary obstacles) positioned on the two sides aligned along the curb-side edge of the road, as shown in Figure 2. The boundaries define the parking space’s length $L_{park}$ and width $W_{park}$. In our setup, other obstacles are located at least $D_{obst}$ away from the parking spots to leave the drivable space. We denote the parking vehicle’s width as $W$, length as $L$, and the distance from the start to the target position as $D_{park}$. Table I shows the categorization based on the aforementioned parameters, with narrower parking lots classified as more difficult. Since parallel parking is more challenging than vertical parking in practice, we introduce an extreme difficulty level for it [38]. Besides, because larger parking distances require more maneuvering over a longer distance and increase the number of obstacles encountered along the way, we rank the scenarios that $D_{park}>15.0$ the complex scenario. Note that we do not specify the initial orientation or position of the vehicle, meaning the starting conditions for the vehicle can be any collision-free configuration, which adds to the difficulty and diversity of scenarios.
@@ V-C2 Further comparison with Hybrid A*
Figure 4: The visualization of the planning process and results of the Hybrid A* (a-1)-(e-1) and the proposed HOPE (a-2)-(e-2). The blue-to-green gradient rectangles represent the states explored during the algorithm’s search process, while the yellow curves indicate the final path planning result. In the normal parallel parking case shown in (a), both methods provide concise path planning results. In the vertical parking scenario shown in (b) and the normal dlp scenario in (d), although both methods succeed in planning, our approach yields more reasonable results. Meanwhile, in the narrow parallel parking scenario (c) and the scenario requiring parking with the front of the vehicle facing inward (e), the Hybrid A* algorithm fails to plan, while our approach succeeds.
@@ V-D1 Hybrid policy with RS-curve
We designed experiments to investigate how the use of the RS method influences the performance of the hybrid policy. A hyperparameter in the hybrid strategy is the threshold distance $d_{\text{rs}}$. The RS policy is considered only when the vehicle is closer to the target position than $d_{\text{rs}}$. The results show that even reducing $d_{\text{rs}}$ to 1 m has less than a 5% impact on the success rate while increasing $d_{\text{rs}}$ does not lead to a significant improvement. This result indicates that while RS curves assist in training the hybrid policy, the reinforcement learning agent does not overly rely on the RS method, allowing it to outperform rule-based approaches.
@@ V-D2 Other proposed modules
More ablation experiment results are presented in Table V. In the absence of the action mask mechanism, the success rate declines by more than 20% in complex parallel parking scenarios and 30% in extreme scenarios. This significant drop in performance indicates that utilizing an action mask to influence and constrain the agent’s action is crucial for training effectiveness. Besides, replacing the transformer with a multi-layer perceptron leads to a decline of more than 10% in several scenarios, demonstrating the importance of multi-modal information fusion. Excluding the action mask or BEV images as inputs to the neural network also impacts the agent’s performance. Using the auto-encoder to pre-train the image encoder can enhance the training for the PPO-based agent but has a relatively minor impact on the SAC-based agent.
@@
-AMP 96.5 49.1 96.3 36.5 23.4 78.2 57.2
MLP 99.8 90.6 98.8 82.3 72.1 97.8 87.9
-AMI 99.9 96.4 98.7 91.9 84.9 97.2 92.3
-BEV 99.8 97.8 98.7 97.9 97.6 98.1 96.5
-AE 100.0 99.8 100.0 99.0 96.7 99.7 99.0
-AMP: omit the action mask mechanism in post-processing. MLP: use MLP instead of transformer as the network backbone. -AMI: exclude the use of the action mask in network input. -BEV: exclude the BEV image in network input. -AE: not employ auto-encoder to pre-train the image encoder.
@@ V-E Real-World Experiments
We also conducted real-world experiments in an indoor parking garage to show the applicability of our method in the real world. Our experimental platform was a Changan CS55E-Rock vehicle with a dimension of $4.6m\times 1.9m$ and we fine-tuned our model with the vehicle’s parameter in the same simulation environment. More details about the experimental platform and the overall system can be found in the appendix E.
@@
Figure 6: The illustration of our real-world experiments in scenarios of vertical parking (a-1,2), parallel parking (b-1,2), and deadend parking (c-1,2). The initial position is shown in the left images and the right images show the target parking lots where the vehicle is positioned.
@@ VI Conclusion and Future Work
To better train and evaluate the proposed method, we introduce a difficulty ranking approach for parking scenarios. The difficulty ranking standard categorizes scenarios into normal, complex, and extreme based on obstacles and parking space conditions. We utilized a simulator and a real-world parking lot dataset to construct training and testing scenarios. Comparisons were made against other rule-based and RL-based methods. Comprehensive experiments were designed to illustrate the effectiveness of the proposed method. The results indicate that our approach outperforms the Hybrid A* algorithm in planning success rates across all types of scenarios, particularly in complex and extreme situations. Compared to the typical approach of directly using reinforcement learning for training, our proposed method combines the advantages of both rule-based and learning-based methods, resulting in an effectively trained planner with generalization capabilities across different scenarios.
@@
This paper shows that the learning-based approach can serve as a promising tool in dealing with complex and diverse parking scenarios. Our work will be open-sourced to support the following research. The proposed HOPE serves as a potential planner to deal with the intricate scenarios where the rule-based methods fail to plan a feasible path. Besides, since only static scenarios are considered in our work, our method can be integrated with a dynamic obstacle avoidance system to handle real-world participants to ensure compatibility and safety in dynamic environments. In the future, more efforts could be made to apply the learning-based method directly in parking scenarios with dynamic obstacles. To support this, enhancing simulation environments with interactive traffic participant models that realistically react to the autonomous vehicle would greatly improve the training and evaluation of such methods.
Appendix A Implementation of information fusion transformer
In our approach, both vectorized and visual representations are employed to convey state information. These inputs of the network include:
  • Obstacle distance (vector-based) $l_{t}$: the value of the nearest distance to obstacles at certain angles, as defined in section IV-C1, which is similar to the scan results of a single-line lidar. The size of this vector is fixed at 120 in practice, with each element representing the minimal obstacle distance within an angular range of $\pi/60$ radians.
  • Action mask (vector-based) $f_{am}$: a discretization of action mask that describes the max valid step size at different steering angles, as defined in section IV-C2. The size of this vector is 42, with 21 elements representing the mask values for forward planning and the remaining 21 for backward planning.
  • Bird-eye-view information (image-based) $I_{BEV}$: a low-resolution depiction of the drivable area and target parking spot. The historical trajectory of the ego vehicle is rendered using a gradient color scheme. The image size is $H_{img}\cdot W_{img}$.
Figure 7: The success frequency of different RS curve types

### v3

@@ Abstract
Automated parking stands as a highly anticipated application of autonomous driving technology. However, existing path planning methodologies fall short of addressing this need due to their incapability to handle the diverse and complex parking scenarios in reality. While non-learning methods provide reliable planning results, they are vulnerable to intricate occasions, whereas learning-based ones are good at exploration but unstable in converging to feasible solutions. To leverage the strengths of both approaches, we introduce Hybrid pOlicy Path plannEr (HOPE). This novel solution integrates a reinforcement learning agent with Reeds-Shepp curves, enabling effective planning across diverse scenarios. HOPE guides the exploration of the reinforcement learning agent by applying an action mask mechanism and employs a transformer to integrate the perceived environmental information with the mask. To facilitate the training and evaluation of the proposed planner, we propose a criterion for categorizing the difficulty level of parking scenarios based on space and obstacle distribution. Experimental results demonstrate that our approach outperforms typical rule-based algorithms and traditional reinforcement learning methods, showing higher planning success rates and generalization across various scenarios. We also conduct real-world experiments to verify the practicability of HOPE. The code for our solution is openly available on https://github.com/jiamiya/HOPE.
@@ I Introduction
Automated parking is a tempting technology to improve driving safety and efficiency [1]. An automated parking system comprises several vital components, including perception, planning, and control, in which path-planning algorithms play a crucial role [2]. In parking scenarios, the path-planning task involves generating a feasible path from the start position to the target parking spots under specific physics constraints. Compared to other scenarios, finding a feasible path in parking scenarios is usually more challenging because of the lower error tolerance of the target spot and the lack of navigational reference lines [3]. Additionally, the limited space and surrounding obstacles shrink the number of potential solutions. While existing path-planning approaches have proved practical in most simple scenarios, their inherent difficulty in understanding the surrounding environment may lead to planning failures, especially when scenarios become more intricate [4].
Learning-based planners have the potential to understand environments and intelligently plan routes through a data-driven approach, diverging from reliance on pre-defined planning methods rooted in human priors. While expert data is used as the ground truth for imitation learning in various tasks, the scarcity of large-scale datasets for parking scenarios necessitates researchers to collect data manually [5]. This supervised learning approach risks overfitting the model to specific parking strategies due to the insufficient diversity in the training scenarios. Meanwhile, Reinforcement Learning (RL) has gathered increasing attention in the field of autonomous driving [6]. Indeed, through interaction with the environment, RL methods enable the training of agents without labeling trajectory ground truth. Nonetheless, training RL agents in complex and diverse scenarios remains a challenging task [7]. It is straightforward for the agent to fit into fixed parking strategies but poses greater difficulty in obtaining generalization capability across various parking scenarios. Moreover, the agent faces challenges in exploring effectively, especially in complex scenarios with narrow parking spaces, significantly impacting the training efficiency.
@@
  • We develop a hybrid policy method for parking path planning which achieves over 97% success rate across diverse and challenging parking scenarios and verifies its generalization ability with real-world experiments.
  • We propose to implement an action mask in the path planning task. This mechanism excludes improper actions for reinforcement learning agents and significantly improves training efficiency and performance.
@@ III-A Reinforcement learning for path planning
The RL problem can be addressed as policy learning in a Markov decision process (MDP) defined by a 4-element tuple $(\mathcal{S},\mathcal{A},p,r)$, where $\mathcal{S}$ is the state space, and $\mathcal{A}$ is the action space. In the path-planning problem, $s_{t}\in\mathcal{S}$ includes the vehicle position and orientation $p_{t}=(x_{t},y_{t},\theta_{t})$ as well as other observable information about obstacles and target parking space. We selected velocity $v$ and steering angle $\delta$ as the action space $\mathcal{A}=\{a=(v,\delta)\}$. $p$ denotes the state transition probability density of the next state $s_{t+1}$ given the current state $s_{t}$ and action $a_{t}$. In practice, the transition is modeled using the single-track bicycle model, as implemented in [32], and this uncertainty is not considered in a deterministic environment. A reward $r=r(s_{t+1},a_{t})$ is given by the environment $E$ based on the state and action in each interaction step. The objective is to learn the policy $\pi(a_{t}|s_{t})$ that maximizes the future reward $R_{t}=\sum_{i=t}^{T}\gamma^{(i-t)}r(s_{i},a_{i})$ with a discounting factor $\gamma\in[0,1]$. Given the start point $p_{0}$ and the target $p_{T}$, the feasible parking path $P=\{p_{0},p_{1},p_{2},...,p_{T}\}$ can be obtained iteratively using the policy $\pi$.
@@
To demonstrate the improvement of our proposed method across different reinforcement learning approaches, in this paper, we chose the commonly used on-policy algorithm, Proximal Policy Gradient (PPO) [33], and off-policy algorithm, Soft Actor-Critic (SAC) [34], to obtain the reinforcement learning policy $\pi_{\theta}$ in HOPE.
@@ IV-A Overview of architecture
The overall RL-based path planning framework and network structure are shown in Figure 1. To enhance training efficiency, we employ a hybrid policy reinforcement learning approach. Our proposed HOPE combines the learnable policy from original reinforcement learning with a rule-based policy derived from the RS curve. In each interaction step, the agent outputs the action, namely the single-step path planning result, based on the current state given by the environment. The action is then adjusted using the action mask mechanism before interacting with the environment. We employ four forms of network input as state representation, including:
  • Obstacle distance (vector-based) $l_{t}$: the value of the nearest distance to obstacles at certain angles, which will be introduced later in section IV-C1.
  • Action mask (vector-based) $f_{am}$: the representation of max valid step size at different steering angles, which will be introduced in section IV-C2.
  • Bird-eye-view information (image-based) $I_{BEV}$: a low-resolution depiction of the drivable area, target parking spot, and the historical trajectory of the ego vehicle.
A transformer-based structure with learnable view encoding is designed to fuse the inputs and get the outputs in the actor and critic network [35]. We also utilize the auto-encoder structure to pre-train the image encoder. More details of our network and reward function are discussed in the appendix A and B. The pseudo-code of our method is shown in Algorithm 1.
@@ IV-B1 Rule-based policy from Reeds-Shepp curve
  • We calculate the shortest K curves and assess their feasibility in ascending order of length, selecting the shortest feasible, collision-free path as the final route.
  • We include paths of the $S(|)C(|)S$ form. These paths are excluded in the original RS method because they are strictly suboptimal in obstacle-free spaces. However, in environments with obstacles, these paths may still be feasible and beneficial, as indicated in Appendix D.
@@ IV-C1 Efficient estimation of action mask
While solving for the optimal $l_{ij}$ considering only obstacle point at $i^{th}$ angle $\omega_{i}$ and vehicle steering angle $\delta_{j}$ is equivalent to obtaining the maximum velocity, denoted as $v^{*}_{ij}=\mathcal{E}_{ij}^{-1}(l_{t}[i])$, calculating all envelope distances corresponding to each $i$ and $j$ in each interaction is computationally expensive, and the inverse of $\mathcal{E}_{ij}$ may not always exist. To deal with this issue, we propose the pre-calculation of the anchoring distance $\hat{l}_{ij}$ for $K_{v}$ discretized velocities:
$\hat{l}_{ij}[k]=\mathcal{E}_{ij}(\hat{v}[k]),\ \hat{v}[k]=v_{max}\frac{k}{K},\ k=0,1,...K_{v}.$ (12)
@@ V-A Scenario difficulty ranking
To better train and evaluate our approach, we rank the difficulty for parking scenarios into normal, complex, and extreme levels based on existing standards ISO 20900 and GB/T 41630-2022 [37, 38]. Generally, a parking space is defined by two boundary parking vehicles (or boundary obstacles) positioned on the two sides aligned along the curb-side edge of the road, as shown in Figure 2. The boundaries define the parking space’s length $L_{park}$ and width $W_{park}$. In our setup, other obstacles are located at least $D_{obst}$ away from the parking spots to leave the drivable space. We denote the parking vehicle’s width as $W$, length as $L$, and the distance from the start to the target position as $D_{park}$. Table I shows the categorization based on the aforementioned parameters, with narrower parking lots classified as more difficult. Since parallel parking is more challenging than vertical parking in practice, we introduce an extreme difficulty level for it [38]. Besides, since larger parking distances require more maneuvering over a longer distance and increase the number of obstacles encountered along the way, we rank the scenarios that $D_{park}>15.0$ the complex scenario. Note that we do not specify the initial orientation or position of the vehicle, meaning the starting conditions for the vehicle can be any collision-free configuration, which adds to the difficulty and diversity of scenarios.
@@ V-C2 Further comparison with Hybrid A*
Figure 4: The visualization of the planning process and results of the Hybrid A* (a-1)-(e-1) and the proposed HOPE (a-2)-(e-2). The blue-to-green gradient rectangles represent the states explored during the algorithm’s search process, while the yellow curves indicate the path planning result. In the normal parallel parking case shown in (a), both methods provide concise path planning results. In the vertical parking scenario shown in (b) and the normal dlp scenario in (d), although both methods succeed in planning, our approach yields more reasonable results. In the narrow parallel parking scenario (c) and the scenario requiring parking with the front of the vehicle facing inward (e), the Hybrid A* fails to plan, while our approach succeeds.
@@ V-D1 Hybrid policy with RS-curve
We designed experiments to investigate how the RS method usageinfluences the performance of the hybrid policy. A hyperparameter in the hybrid strategy is the threshold distance $d_{\text{rs}}$. The RS policy is considered only when the vehicle is closer to the target position than $d_{\text{rs}}$. The results show that even reducing $d_{\text{rs}}$ to 1 m has less than a 5% impact on the success rate while increasing $d_{\text{rs}}$ does not lead to a significant improvement. This result indicates that while RS curves assist in training the hybrid policy, the reinforcement learning agent does not overly rely on the RS method, allowing it to outperform rule-based approaches.
@@ V-D2 Other proposed modules
More ablation experiment results are presented in Table V. In the absence of the action mask mechanism, the success rate declines by more than 20% in complex parallel parking scenarios and 30% in extreme scenarios. This significant drop in performance indicates that utilizing an action mask to influence and constrain the agent’s action is crucial for training effectiveness. Besides, replacing the transformer with a multi-layer perceptron leads to a decline of more than 10% in several scenarios, demonstrating the importance of multi-modal information fusion. In situations where BEV data is unavailable, our model can still work effectively by removing the BEV-related input branch from the transformer module. This adjustment does not lead to significant performance drops, and in some scenarios, it even achieves comparable performance to the original. Using the auto-encoder to pre-train the image encoder can enhance the training for the PPO-based agent but has a relatively minor impact on the SAC-based agent.
@@
W/O AMP 96.5 49.1 96.3 36.5 23.4 78.2 57.2
W/O TF 99.8 90.6 98.8 82.3 72.1 97.8 87.9
W/O AMI 99.9 96.4 98.7 91.9 84.9 97.2 92.3
W/O BEV 99.8 97.8 98.7 97.9 97.6 98.1 96.5
W/O AE 100.0 99.8 100.0 99.0 96.7 99.7 99.0
W/O AMP: without the action mask in post-processing. W/O TF: use MLP instead of transformer as the backbone. W/O AMI: without the action mask in network input. W/O BEV: without the BEV image in network input. W/O AE: not employ auto-encoder to pre-train the image encoder.
@@ V-E Real-World Experiments
We also conducted real-world experiments in an indoor parking garage to show the applicability of our method in the real world. Our experimental platform was a Changan CS55E-Rock vehicle with a dimension of $4.6m\times 1.9m$ and we fine-tuned our model with the vehicle’s parameter in the same simulation environment. More details about the platform and the overall system can be found in the appendix C.
@@
Figure 6: The illustration of our real-world experiments in scenarios of vertical parking (a-1,2), parallel parking (b-1,2), and dead-end parking (c-1,2). The initial position is shown in the left images and the right images show the target parking lots where the vehicle is positioned.
@@ VI Conclusion and Future Work
To better train and evaluate the proposed method, we introduce a difficulty ranking approach for parking scenarios, which categorizes scenarios into normal, complex, and extreme based on obstacles and parking space conditions. We utilized a simulator and a real-world dataset to construct training and testing scenarios. Comprehensive experiments show the effectiveness of our method against other rule-based and RL-based methods. The results indicate that our approach outperforms the Hybrid A* in planning success rates across all types of scenarios, particularly in complex and extreme situations. Compared to the typical approach of directly using reinforcement learning, our proposed method combines the advantages of both rule-based and learning-based methods, resulting in an effective planner with generalization capabilities across different scenarios.
@@
This paper shows that the learning-based approach can serve as a promising tool in dealing with complex and diverse parking scenarios. The proposed HOPE serves as a potential planner to deal with the intricate scenarios where the rule-based methods fail to plan a feasible path. Besides, since only static scenarios are considered in our work, our method can be integrated with a dynamic obstacle avoidance system to handle real-world participants to ensure compatibility and safety in dynamic environments. In the future, more efforts could be made to apply the learning-based method directly in parking scenarios with dynamic obstacles. To support this, enhancing simulation environments with interactive traffic participant models that realistically react to the autonomous vehicle would greatly improve the training and evaluation of such methods.
Appendix A Implementation detials